part 1 Unit Cost, Valuing Goods Transferred Out and EWIP During April, the grinding department of...

part 1

Unit Cost, Valuing Goods Transferred Out and EWIP

During April, the grinding department of Tranx Inc. completed and transferred out 312,000 units. At the end of April, there were 156,000 units in process, 60% complete. Tranx incurred manufacturing costs totaling $2,433,600.

Required:

1. Calculate the unit cost.   $per unit

2. Calculate the cost of goods transferred out and the cost of ending WIP.

Cost of goods transferred out $
Cost of ending WIP $
Answer #1

Ans:

1. Unit Cost= Total Cost/Total Units

Total Units:

=> Total transferred= 312,000*100%= 312,000 units

Total Ending WIP= 156,000*60% = 93,600 Units

Total Units= 312,000+93,600

=> 405,600 units

Total Cost= $2,433,600

Unit Cost per Unit= $2,433,600/405,600

=> $6 per unit

2. Cost of Goods transferred: 312,000 units* $6

=> $1,872,000

Cost of Ending WIP: 93,600 units*$6

=> $561,600